OpenStudy (anonymous):

150 days to a simplified fraction of a year

OpenStudy (anonymous):

Do you mean 150 over 365? 150 --- 365 simplified it would be 30 over 75 30 -- 75

OpenStudy (anonymous):

no , the question is telling me to assume a 360 year day as needed and its saying "150 days to a simplified fraction of a year" im not sure how to solve this but the answer should be a fraction.

OpenStudy (anonymous):

number 5

OpenStudy (here_to_help15):

its C

OpenStudy (here_to_help15):

150/360=15/36=5/12

OpenStudy (anonymous):

great . thank you. so how do i solve number 7?

OpenStudy (here_to_help15):

um let me check and can u fan me?

OpenStudy (here_to_help15):

and medal

OpenStudy (anonymous):

got you

OpenStudy (anonymous):

can someone solve # 17 for me please